# Read Between the Layers: Leveraging Intra-Layer Representations for Rehearsal-Free Continual Learning with Pre-Trained Models

### v1

@@ 5.5 Analysis of intra-layer representation depth
To provide more insights into the impact of the intra-layer representation depth on LayUP performance, we plot in Tab. 3 the average accuracy after training for different choices of $k$ using a Vit-B/16-IN1K backbone and AdaptFormer [5] as PETL method.
$k$ CIFAR IN-R IN-A CUB OB VTAB Cars
[…]
Table 3: Average accuracy (%) after training: LayUP performance for different values of $k$ for a pre-trained Vit-B/16-IN1K and FSA with AdaptFormer. The 1st, 2nd, and 3rd highest scores and the 1st, 2nd, and 3rd lowest scores are highlighted.
We observe that LayUP configurations with $k\geq 4$ consistently surpass the final layer classifier (i.e., $k=1$), which shows that class-prototype methods for CL can generally benefit from intra-layer representations. Although the optimal choice of $k$ differs for each dataset, the results are indicative of a choice of larger $k$. As a larger value of $k$ directly corresponds to increased memory and runtime complexity during inference, its choice should be carefully weighed up, taking into account demands and restrictions at different dimensions.
@@ 5.6 How universal are intra-layer prototypes?
We are interested in determining how universal the intra-layer prototypes are for classification, i.e., how broad the range of classes is that benefit from LayUP. For this purpose, we run phase B of Alg. 1 and provide the fraction of classes for each dataset that has an increased or equal final accuracy score with LayUP for $k=6$ compared with LayUP for $k=1$ (last layer Gram-inverted classifier, cf. Eq. 2).
CIFAR IN-R IN-A CUB OB VTAB Cars
$94$% $79$% $81$% $89$% $80$% $94$% $83$%
Table 4: Percentage of classes per dataset that are better or equally classified with LayUP for $k=6$ compared with the last layer Gram-inverted classifier (Eq. 2).
Tab. 4 shows that between 80% and 94% of the total classes in each dataset are better or equally classified upon introducing intermediate representations to class-prototype construction, indicating the universal applicability of intra-layer prototypes to a broad range of tasks, classes, or domains.
@@ 6 Conclusion
In this paper, we propose LayUP, a simple yet effective rehearsal-free class-prototype method for continual learning with pre-trained models. LayUP leverages intra-layer representations of a pre-trained feature extractor to increase robustness and generalizability, especially under large domain gaps and in low-data regimes. It further computes second-order feature statistics to decorrelate class prototypes, combined with parameter-efficient first-task adaptation. Extensive experiments across a range of image classification datasets and continual learning settings demonstrate that LayUP performs strongly against competitive baselines at a small fraction of their memory and compute cost. In future work, we will further investigate integrating class-prototype methods with continual adaptation of pre-trained models beyond first-task adaptation, which is particularly beneficial for learning under multiple distributional shifts.
Supplementary Material
@@ A.1 Optimization of ridge regression parameter $\lambda$
We perform the optimization over the regression parameter $\lambda$ (cf. Alg. 1) as follows: For every task $t$, we perform a random 80:20 split of the training data stratified by ground truth labels. We then update ${\bm{G}}$ and ${\bm{c}}_{y}\ \forall\ y\in\mathcal{Y}_{t}$ (which were already updated during task $11,\dots,t-$ 1 training) for the first 80% of training data of task $t$ and then choose the value of $\lambda\in\{10^{-3},10^{-2},10^{-1},10^{0},10^{1},10^{2},10^{3}\}$ that yields highest per-sample accuracy on the remaining 20% of training data. Finally, we update ${\bm{G}}$ and ${\bm{c}}_{y}\ \forall\ y\in\mathcal{Y}_{t}$ for those 20% of training data of task $t$ and repeat the same process with $t+1$.
@@ A.2 Data augmentation
During the training process, data augmentation was applied to all datasets, incorporating random cropping of the images to varying sizes, ranging from 70% to 100% of their original dimensions, while maintaining an aspect ratio between 3:4 and 4:3. After the resizing, images were randomly flipped horizontally and brightness, contrast, saturation, and hue were varied randomly in a 10% range. Finally, the images were center-cropped to 224×224 pixels for all datasets, except for CIFAR-100, where images were directly resized from the original 32×32 to 224×224 pixels. During inference, images of all datasets were resized to 224x224 pixels without further modification.
@@ Appendix B Datasets
We provide a summary of the datasets compared in the main experiments in Tab. 5. CIFAR-100 (CIFAR) contains 100 classes of natural images of different domains and topics and can be considered relatively in-distribution with the pre-train domain of ImageNet-1K and ImageNet-21K. ImageNet-R (IN-R) contains image categories overlapping with ImageNet-1K, but is a selection of out-of-distribution samples for the pre-train dataset that are either hard examples or newly collected data of different styles. ImageNet-A (IN-A) likewise has overlapped categories with ImageNet-1K, but comprises real-world adversarially filtered images that fool existing ImageNet pre-trained classifiers. The Caltech-UCSD Birds-200-2011 (CUB) dataset is a specialized collection of labeled images of 200 bird species, encompassing a diverse range of poses and backgrounds. OmniBenchmark (OB) serves as a compact benchmark designed to assess the generalization capabilities of pre-trained models across semantic super-concepts or realms, encompassing images of 300 categories that represent distinct concepts. The Visual Task Adaptation Benchmark (VTAB) as used in our work is a composition of the five datasets Resisc45 [8], DTD [10], Pets [46], EuroSAT [15], and Flowers [42], that are learned consecutively to emulate the emergence of different domains. The Stanford Cars-196 (Cars) dataset comprises images of cars belonging to one of 196 unique combinations of model and make.
Original CL $T$ $N_{\text{train}}$ $N_{\text{val}}$ $C$
CIFAR [29] [51] 10 50 000 10 000 100
IN-R [16] [64] 10 24 000 6 000 200
IN-A [17] [72] 10 6 056 1 419 200
CUB [60] [72] 10 9 465 2 323 200
OB [70] [72] 10 89 668 5 983 300
VTAB [68] [72] 5 1 796 8 619 50
Cars [28] [69] 10 8 144 8 041 196
Table 5: Overview of datasets: Original publication, formulation for the CL setting, number of tasks ($T$) in the main experiments in Sec. 5, number of training samples ($N_{\text{train}}$), number of validation samples ($N_{\text{val}}$), and number of classes ($C$).
@@ Appendix C Additional Results
We use the following two metrics in our work for experimental evaluation: Average accuracy $A_{t}$ (following the definition of [36]) and average forgetting $F_{t}$ (following the definition of [3]). Average accuracy is defined as
@@ C.1 Variability across seeds and performance over time
LayUP accuracy and forgetting for all seven datasets over the course of class-incremental training (Phase B in Alg. 1) is depicted in Fig. 3. There are almost no differences between final results after the last task across seeds, which indicates LayUP being insensitive to different task orders and initialization of PETL parameters. The highest variability can be observed for the VTAB benchmark, where each task represents a completely new domain. Therefore, the task order–and consequently the domain that the representations are adjusted to during first session adaptation–has a greater influence on the generalization capabilities of the model during CL. As expected, the average forgetting increases as more tasks are being introduced, as the model has more classes to tell apart during inference.
@@ C.2 Experiments with different layer choice $k$
To analyze the learning behavior over time in the CIL setting among different choices of maximum representation depth $k$ for prototype construction, we plot average accuracy and forgetting for $k=1$ (last layer only), $k=6$, and $k=12$ (all layers) in Fig. 4. Clearly, the classification performance based on last layer representations only is inferior to intra-layer representations as used in LayUP, as it leads to both a lower accuracy and a higher forgetting rate. At the same time, there is no difference in performance between $k=6$ and $k=12$ evident, indicating that early layers of the model do not add meaningful knowledge to the classification. Given that the choice of $k$ is subject to a trade-off between performance gain and memory and computational cost, the results confirm $k=6$ as used in the main experiments in Sec. 5 to be a reasonable choice.
@@ C.3 Experiments with different backbones and PETL methods
Following prior works [72, 40], we experiment with different PETL methods and ViT-B/16 models for LayUP, as the additional benefit from additional fine-tuning of the representations differs depending on the characteristics of the downstream domain [44]. Results are shown in Fig. 5.
In all but one dataset, we found adapter methods (AdaptFormer) to be superior to be superior to prompt learning (VPT) or feature modulation (SSF). However, there is no combination of PETL method and pre-trained model that generally outperforms all other variants. Such results confirm the findings of [44] and underline the importance of considering different pre-training schemes and strategies for additional fine-tuning.
Figure 5: Average accuracy (left) and average forgetting (right) after training on each task $t$ in the CIL setting: Comparison of different PETL methods (AdaptFormer [5], SSF [34], and VPT [22]) and ViT-B/16 models.
@@ C.4 Experiments with different task counts $T$
To show whether a benefit of intra-layer representations can be observed for different task counts, we compare average accuracy scores after training for six different datasets, three different task counts ($T\in\{5,10,20\}$) and three different choices of $k$ last layers for class-prototype generation. $k=1$ corresponds to classification only based on the last layer (i.e., final) representations of the backbone, as it is done in prior work. $k=6$ means to concatenate layer-wise features from the latter half of the network layers. Finally, $k=12$ uses concatenated features of all layers of the pre-trained ViT for classification. Results are presented in Tab. 6.
Performance scores across task counts and datasets are consistently higher for intra-layer representations ($k=6$ and $k=12$) compared with last-layer-only representations ($k=1$). However, differences in performance between $k=6$ and $k=12$ are not apparent (with the exception of a $T=20$ split on the CUB dataset, which can be attributed to some sensitivity to initialization of adapter parameters). Considering that a higher $k$ corresponds to an increased computational and memory demand as ${\bm{G}}$ and ${\bm{c}}_{y}$ increase in dimension, this confirms $k=6$ as reasonable choice for maximum representation depth.
@@
1 5 89.9 80.6 62.5 85.3 73.2 75.6
10 89.1 79.1 61.4 85.5 71.6 75.6
20 86.3 77.8 59.5 83.5 72.2 75.5
6 5 91.3 83.0 64.5 87.6 78.0 82.2
10 90.8 81.9 62.5 87.7 76.6 81.8
20 90.3 80.4 61.6 69.5 77.8 81.8
12 5 91.6 82.7 62.7 87.1 77.6 81.8
10 91.1 82.1 62.7 87.5 73.8 81.7
20 90.5 80.3 56.8 86.0 78.3 81.9
Table 6: Average accuracy (%) after training: Comparison of different task counts $T$ for $k=1$ (prototype construction from last layer only), $k=6$ (as used in Sec. 5), and $k=12$ (prototype construction from all network layers). Scores listed are for AdaptFormer and ViT-B/16-IN1K. VTAB has a fixed number of datasets that are treated as tasks (thus $T=5$) and is therefore omitted in the comparison.

### v2

@@ 5.5 Analysis of intra-layer representation depth
To provide more insights into the impact of the intra-layer representation depth on LayUP performance, we plot in Tab. 4 the average accuracy after training for different choices of $k$ using a ViT-B/16-IN1K backbone and AdaptFormer as PETL method.
The final layer classifier ($k=1$) exhibits the lowest performance in comparison to all other $k$ values across all datasets, except for the split VTAB dataset. Conversely, LayUP configurations with medium or large $k$ values almost consistently outperform those with small $k$ values, demonstrating the broad advantages of leveraging representations from multiple intermediate layers for class-prototype construction. While the optimal choice of $k$ varies across datasets, performance generally improves or stabilizes with increasing $k$ values, thereby selecting intermediate features from a higher number of consecutive layers. However, a larger $k$ value corresponds to increased memory and runtime complexity during inference, necessitating careful consideration of its choice based on the demands and constraints at different dimensions.
@@ 5.6 How universal are intra-layer prototypes?
We aim to assess the universality of intra-layer prototypes for classification, specifically, the breadth of classes benefiting from LayUP. To this end, we report the proportion of classes for each dataset whose test accuracy scores are higher, equal, or lower with LayUP compared to the final layer ridge classifier (cf. Eq. 2). Beyond merely counting the classes that are better predicted by either method, we also examine the extent of the benefit one method has over the other per improved class. Consequently, we calculate the average difference in accuracy scores between LayUP and the final layer classifier ($k=1$), and vice versa, for each class that shows improvement.
Figure 6: LayUP vs. classification from the final representation layer. Left: Percentage of classes per dataset that are better classified with LayUP (with $k=6$) compared with the last layer ridge classifier (equivalent to LayUP with $k=1$). Right: Difference of absolute accuracy per improved class between LayUP and the last layer classifier ($k=1$), and vice versa.
Fig. 6 illustrates that a higher percentage of classes (between 18% and 72%) is more effectively classified upon introducing intermediate representations to class-prototype construction. Although the difference in the number of improved classes is not as pronounced for IN-A and CUB compared with all other benchmarks, the relative difference in accuracy per improved class between LayUP and the final representation layer classifier reaches as high as 39% for IN-A and 22% for CUB. This indicates that for these two datasets, the introduction of intra-layer representations significantly benefits a few classes rather than slightly benefiting a large number of classes, as observed with the other benchmarks. Despite the variability with respect to the number of classes per dataset that benefit from LayUP and the extent of the benefit, our results demonstrate the universal applicability of intra-layer prototypes across a broad spectrum of tasks, classes, or domains.
@@ 6 Conclusion
In this paper, we propose LayUP, a simple yet effective rehearsal-free class-prototype method for continual learning with pre-trained models. LayUP leverages intra-layer representations of a pre-trained feature extractor to increase robustness and generalizability, especially under large domain gaps and in low-data regimes. It further computes second-order feature statistics to decorrelate class prototypes, combined with parameter-efficient first-task adaptation. Extensive experiments across a range of image classification datasets and continual learning settings demonstrate that LayUP performs strongly against competitive baselines while requiring significantly less memory and computational complexity. In future work, we will further investigate integrating class-prototype methods with continual adaptation of pre-trained models beyond first-task adaptation, which is particularly beneficial for learning under multiple distributional shifts. Intra-layer representations serve as a potent source of knowledge that is inherently available in pre-trained models. Based on the findings made in this work, we encourage further exploration into making explicit use of these representations (i.e., reading between the layers) in future approaches to continual learning.
Appendix
@@ A.1 Optimization of ridge regression parameter $\lambda$
The optimization process for selecting the regression parameter $\lambda$ (cf. Alg. 1) proceeds as follows: For each task $t$, we stratify the training data by ground truth labels and perform a four-fold split. For each split, we temporarily update ${\bm{G}}$ and ${\bm{c}}_{y}$ for all $y\in\mathcal{Y}_{t}$, and calculate the accuracy for values of $\lambda\in\{10^{-8},10^{-4},10^{-3.5},10^{-3},10^{-2.5},10^{-2},10^{-1.5},10^{-1},10^{-0.5},10^{0},10^{0.5},10^{1},10^{1.5},10^{2},10^{2.5},10^{3}\}$ to determine which yields the highest per-sample accuracy on the remaining training data fold. We then calculate the mean accuracy across all folds for each $\lambda$ and select the value that results in the highest overall accuracy. Finally, we update ${\bm{G}}$ and ${\bm{c}}_{y}$ for all $y\in\mathcal{Y}_{t}$ using the complete training dataset for task $t$, and repeat the process for task $t+1$. Note that this procedure, when performed at time $t$, does not require access to data from any prior task, thus maintaining full compatibility with the rehearsal-free CL setting.
@@ A.2 Data augmentation
During CL training, data augmentation is applied to all datasets, incorporating random cropping of the images to varying sizes, ranging from 70% to 100% of their original dimensions, while maintaining an aspect ratio between 3:4 and 4:3. After resizing, images are randomly flipped horizontally, and brightness, contrast, saturation, and hue are varied randomly within a 10% range. Finally, the images are center-cropped to 224×224 pixels for all datasets, except for CIFAR-100, where images are directly resized from the original 32×32 to 224×224 pixels. During inference, images of all datasets are resized to 224x224 pixels without further modification.
@@ Appendix B Datasets
We summarize the datasets compared in the main experiments in Tab. 5.
Datasets for the CIL/OCL settings.
CIFAR-100 (CIFAR) comprises 100 classes of natural images across various domains and topics, aligning relatively closely with the pre-training domains of ImageNet-1K and ImageNet-21K in terms of distribution. ImageNet-R (IN-R) consists of image categories that overlap with ImageNet-1K, yet it features out-of-distribution samples for the pre-training dataset, including challenging examples or newly collected data of various styles. Similarly, ImageNet-A (IN-A) shares categories with ImageNet-1K but encompasses real-world, adversarially filtered images designed to deceive existing classifiers pre-trained on ImageNet. The Caltech-UCSD Birds-200-2011 (CUB) dataset is a specialized collection of labeled images of 200 bird species, encompassing a diverse range of poses and backgrounds. OmniBenchmark (OB) serves as a compact benchmark designed to assess the generalization capabilities of pre-trained models across semantic super-concepts or realms, encompassing images of 300 categories that represent distinct concepts. The Visual Task Adaptation Benchmark (VTAB) as used in our work is a composition of the five datasets Resisc45 (Cheng et al., 2017), DTD (Cimpoi et al., 2014), Pets (Parkhi et al., 2012), EuroSAT (Helber et al., 2019), and Flowers (Nilsback & Zisserman, 2006) and is commonly split into $T=5$ tasks in the CL version. The Stanford Cars-196 (Cars) dataset comprises images of cars belonging to one of 196 unique combinations of model and make.
Datasets for the DIL setting.
The Continual Deepfake Detection Benchmark Hard (CDDB-H) is a framework designed to evaluate the performance of deepfake detection models under continuously changing attack scenarios (gaugan, biggan, wild, whichfaceisreal, and san). DomainNet is a multi-source domain adaptation benchmark, comprising the six image style domains real, quickdraw, painting, sketch, infograph, and clipart. As the original training dataset of DomainNet comprises more than 400 000 samples, we use a sub-sampled version of DomainNet (S-DomainNet), that caps the number of training samples at 10 per class and domain, while maintaining the original validation set unchanged. Finally, with the domain-incremental split of ImageNet-R (IN-R (D)), a sequence of fifteen image style domains is learned (sketch, art, cartoon, deviantart, embroidery, graffiti, graphic, misc, origami, painting, sculpture, sticker, tattoo, toy, and videogame).
Setting Name Original CL Formulation $T$ $N_{\text{train}}$ $N_{\text{val}}$ $C$
CIL/OCL CIFAR Krizhevsky (2009) Rebuffi et al. (2017b) 10 50 000 10 000 100
IN-R Hendrycks et al. (2021a) Wang et al. (2022b) 10 24 000 6 000 200
IN-A Hendrycks et al. (2021b) Zhou et al. (2023b) 10 6 056 1 419 200
CUB Wah et al. (2011) Zhou et al. (2023b) 10 9 465 2 323 200
OB Zhang et al. (2022) Zhou et al. (2023b) 10 89 668 5 983 300
VTAB Zhai et al. (2020) Zhou et al. (2023b) 5 1 796 8 619 50
Cars Krause et al. (2013) Zhang et al. (2023) 10 8 144 8 041 196
DIL CDDB-H Li et al. (2023) Li et al. (2023) 5 16 068 10 706 2
S-DomainNet Peng et al. (2019) Peng et al. (2019) 6 20 624 176 743 345
IN-R (D) Hendrycks et al. (2021a) 15 24 000 6 000 200
Table 5: Overview of datasets. Original publication, split formulation for CL setting, number of tasks ($T$) in the main experiments in Sec. 5, number of training samples ($N_{\text{train}}$), number of validation samples ($N_{\text{val}}$), and number of classes ($C$). Although IN-R was originally proposed for domain generalization tasks, we are not aware of any prior work to use it in the DIL setting.
@@ Appendix C Additional Results
We use the following two metrics in our work for experimental evaluation: Average accuracy $A_{t}$ (following the definition of Lopez-Paz & Ranzato (2017)) and average forgetting $F_{t}$ (following the definition of Chaudhry et al. (2018)). Average accuracy is defined as
@@ C.1 Variability across seeds and performance over time
As the choice of the random seed has an influence on both the task order and PETL parameter initialization, we calculate average accuracy and forgetting with standard error for all seven datasets over the course of class-incremental training (phase B in Alg. 1) using seeds 1993-1997. Results are shown in Fig. 7. There are almost no differences between final performance after the last task across seeds, which indicates LayUP being robust to different task orders and initialization of PETL parameters. The highest variability can be observed for the VTAB benchmark, where we have $T=5$ and classes belong to very distinct domains. Therefore, the task order–and consequently the set of classes that the representations are adjusted to during first session adaptation–has a greater influence on the generalization capabilities of the model during CL. As expected, the average forgetting increases as more tasks are being introduced, as the model has more classes to tell apart during inference.
@@ C.2 Experiments with different layer choice $k$
To analyze the learning behavior over time in the CIL setting among different choices of maximum representation depth $k$ for prototype construction, we plot average accuracy and forgetting for $k=1$ (last layer only), $k=6$, and $k=12$ (all layers) in Fig. 8. Clearly, the classification performance based on last layer representations only is inferior to intra-layer representations as used in LayUP, as it yields both a lower accuracy and a higher forgetting rate. At the same time, there is no performance difference between $k=6$ and $k=12$ evident, indicating that the model’s early layers do not contribute meaningful knowledge to the classification process, yet they do not detriment it either. Considering that the choice of $k$ is subject to a trade-off between performance gain and both memory and computational costs, the results corroborate $k=6$, as used in the main experiments detailed in Sec. 5, to be a reasonable choice.
@@ C.3 Experiments with different backbones and PETL methods
Following prior works (Zhou et al., 2023b; McDonnell et al., 2023), we experiment with different PETL methods and ViT-B/16 models for LayUP, as the additional benefit from additional fine-tuning of the representations differs depending on the characteristics of the downstream domain (Panos et al., 2023). Results are shown in Fig. 9.
In all but one dataset, we found adapter methods (AdaptFormer) to be superior to be superior to prompt learning (VPT) or feature modulation (SSF). However, there is no combination of PETL method and pre-trained model that generally outperforms all other variants. Such results confirm the findings of (Panos et al., 2023) and underline the importance of considering different pre-training schemes and strategies for additional fine-tuning.
Figure 9: Average accuracy (left) and average forgetting (right) after training on each task $t$ in the CIL setting: Comparison of different PETL methods (AdaptFormer (Chen et al., 2022), SSF (Lian et al., 2022), and VPT (Jia et al., 2022)) and ViT-B/16 models.
@@ C.4 Experiments with different task counts ${T}$
To show whether a benefit of intra-layer representations can be observed for different task counts, we compare average accuracy scores after training for six different datasets, three different task counts $T\in\{5,10,20\}$, and three different choices of $k$ last layers for class prototype generation. $k=1$ corresponds to classification only based on the last layer (i.e., final) representations of the backbone, as it is done in prior work. $k=6$ means to concatenate layer-wise features from the latter half of the network layers. Finally, $k=12$ uses concatenated features of all layers of the pre-trained ViT for classification. Results are presented in Tab. 6.
Performance scores across task counts and datasets are consistently higher for intra-layer representations ($k=6$ and $k=12$) compared with only final representations ($k=1$). However, there is no considerable difference in performance between $k=6$ and $k=12$. Thus, incorporating intermediate representations from layers in the second half of the model provides sufficient information for class separation, while features from early layers contribute minimally. This supports the observations noted in the preliminary analysis in Sec. 4.1.
@@
1 5 89.5 80.4 61.9 85.0 72.4 75.1
10 89.2 79.2 60.8 85.3 71.6 75.5
20 88.4 77.3 57.1 83.6 72.2 75.6
6 5 91.8 82.8 64.0 87.1 77.4 81.7
10 91.0 81.2 62.2 87.3 77.5 82.5
20 88.8 79.9 59.6 85.9 77.4 82.2
12 5 91.6 83.2 64.0 87.4 78.3 82.1
10 90.8 80.8 62.8 88.0 77.8 82.3
20 88.9 80.2 60.0 86.0 78.7 82.8
Table 6: Average accuracy (%) after training: Comparison of different task counts $T$ for $k=1$ (prototype construction from last layer only), $k=6$ (prototype construction from second half of the network layers, as utilized in Sec. 5), and $k=12$ (prototype construction from all network layers). Scores listed are for AdaptFormer and ViT-B/16-IN1K. VTAB is omitted from the comparison due to its fixed number of datasets from which tasks are constructed ($T=5$).